OpenStudy (anonymous):

how (step-by-step wise) do you calculate parametric representations of a function if given two points? ex. line segment (0,0) to (2,0) and (2,0) to (3,2)

OpenStudy (anonymous):

do you want a separate parametrization for each line segment?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

\[x(t)=t, y(t)=0, \operatorname {for} 0\leq t\leq 2\]

OpenStudy (anonymous):

how did you get that? I know you use the equation r(t) = (1-t)ro - tr1 but when i try that I get x= 2t and y=0 then for the next one i get x=2+t and y= 2 but those are wrong

OpenStudy (anonymous):

your first one is correct the second should be \[x(t)=2+t, y(t)=2t, \operatorname{for} 0\leq t\leq 1\]

OpenStudy (anonymous):

parameterizations are not unique as long as the parametric representations trace out those line segments , you are good

OpenStudy (anonymous):

so then when i evaluate the line integral of integral of (xydx + (x-y)dy) where C consists of those line segments, i can use those parametric equations?

OpenStudy (anonymous):

you will have two integrals, but yes, those should work

OpenStudy (anonymous):

ok, I'll try that. Can i verify my answer with you in a bit?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

great, thank you

OpenStudy (anonymous):

i get 17/3 as my solution

OpenStudy (anonymous):

i got 10 but let me check my work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i reworked the problem and got 10 again

OpenStudy (anonymous):

hmm, for the c1 i get the integral to equal 0 so then the c2 second integral i got 17/3

OpenStudy (anonymous):

I got 2 for c1 and 8 for c2

OpenStudy (anonymous):

hmm ok so my c1 is integral from 0 to 1 (2t)(0)(2)dt + (2t-0)(0)dt which is zero. since my parametric equations are x = 2t and y = 0

OpenStudy (anonymous):

yeah your right, my mistake,

OpenStudy (anonymous):

good catch

OpenStudy (anonymous):

what was your second?

OpenStudy (anonymous):

phew ok good, i thought i really didn't know what i was doing anymore.

OpenStudy (anonymous):

I still get 8

OpenStudy (anonymous):

for my second my parametric equations are x(t) = 2+t and y(t) = 2t

OpenStudy (anonymous):

so its integral from 0 to 1 again of (2+t)(2t)(1)dt + (2+t-2t)(2)dt

OpenStudy (anonymous):

you know what, you are right, I am trying to do too much in my head...

OpenStudy (anonymous):

Ah, yeah i gave up trying to do math in my head since I never get it right.

OpenStudy (anonymous):

said thing is I'll be teaching this stuff next week...ha! thanks for the practice

OpenStudy (anonymous):

oh no way? thats pretty neat. Thank you for helping me with my parameterization confusion.

OpenStudy (anonymous):

17/3 is correct

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

my students know I seldom add correctly....good luck

OpenStudy (anonymous):

Ah yes, it is frustrating when one can do the entire problem and in the end mess it up because of addition. Thank you, good luck to you when teaching it next week!